Jefferson Lab > Physics > User Liaison
Privacy and Security Notice

Users Group: Board of Directors Minutes

Minutes of the meeting on June 11, 2003

BOD Members Present:
Paul Stoler (Chair), Jian-Ping Chen, Ron Gilman, Sebastian Kuhn, Sabine Jeschonnek, Alan Nathan (Past Chair), David Lawrence, Pete Markowitz, Clara Purdue, Doug Higginbotham, Curtis Meyer, Maurizio Ungaro, Mark Pitt, David Armstrong.

Jefferson Lab Representatives Present:
Swapan Chattopadhyay, Dennis Skopik, Kees de Jager, Christoph Leemann, Larry Cardman, Andrew Hutton, Volker Burkert, Rolf Ent.

Introduction of New Members: Paul Stoler

New board members Ron Gilman, Jian-Ping Chen and David Armstrong were introduced and welcomed. Their responsibilities will be Ron: computing issues, Jian-Ping: running experiments, David: PAC issues. Alan Nathan will be in charge of public relations and Sabine Jeschonnek liaison between theory and experiment. Departing members Doug Higginbotham, Sebastian Kuhn and Curtis Meyer were thanked for their contributions. Armstrong was drafted to take the minutes.

Budget: Christoph Leemann

Larry Cardman: 12 GeV upgrade

The PCDR (preliminary conceptual design report) draft is on the WWW, and is open for comments from users until July 3rd. The aim is to have the final document complete by late July. The PCDR is in very good shape, and would form a very good basis for an eventual CDR. Users are thanked for their efforts on the PCDR.

Andrew Hutton: Operations

Christoph Leemann: Chief Scientist Position

An offer was extended two weeks ago to an eminent scientist. The candidate is expected to visit the lab soon, probably in July.

All: User Role in Furthering JLab aspirations

Hall Leaders Reports

Kees de Jager: Hall A

A major source of concern in Hall A has been the septum magnets. The first was delivered to JLab in February. After installation and initial cooldown it was found that the coils didn't adequately cool, apparently there was insufficient direct cooling to the coils. Two temporary fixes were installed, one being direct cooling of the heat shield with liquid He, and the second being direct cooling of the magnet bore with liquid nitrogen. These together allowed the coils to cool enough to go superconducting. After a lengthy training process the magnet reached full current. However, initial commissioning data with beam were baffling. Eventually it was determined that he coils were miswired, and the field was consequently not a dipole. Nevertheless, the forward-angle GDH was able take its low-energy data set with the miswired septum. The magnet was subsequently warmed, and last weekend the miswiring was confirmed and was corrected June 9th. The close up for re-cooling has been started.

The forward-angle GDH experiment will start again mid-July, and should be complete by the down time at the end of August. The vendor has allowed substantially increased JLab involvement with the assembly of the second septum. It is expected to be delivered to JLab in late August. The septum difficulties have forced significant changes to the Hall A schedule. The schedule now has during the Fall the installation of the second septum, repairs to the cooling problems for the first, and the installation of the waterfall target, with the hypernuclear experiment starting Dec. 4th. Following this, the septa will be removed and the BigBite spectrometer will be installed to run the short-range nucleon correlations experiment. The present backlog of experiments in Hall A is about 5 years, with an `off-the-cuff' estimate of the effect of the septum delays being an increase of 3-6 months in the backlog.

Volker Burkert: Hall B

The February - April period was devoted to drift chamber (DC) repairs. A problem with increasing failures in on-board DC electronics (pre-amps) in certain areas of the region III DCs caused an early end to the previous e1e running period. The problem was traced to corrosion on the printed circuit boards, due to poor cleaning at one stage of the manufacturing. Originally the plan was to replace the PC boards on 4 of the segments, but in fact 6 segments were repaired. Subsequently, some channels on segments that were not replaced have been observed to die, and so at the next down all remaining segments will be replaced. In addition the high voltage supplies which were grouped in 96 channels were split to be grouped in 48 channels, so that if one channel fails, only 47 others have to be powered off.

The e1 run group is now taking data, and will end (e1f and e1g completed) at the end of July. This should complete the entire e1 run group of experiments. Then eg2 will start (high energy elecrons on nuclear targets to study hadronization in nuclear matter) followed by a short test for DVCS in preparation for the 6 GeV DVCS run planned for next year, and eg2 until the end of the year.

The backlog in Hall B is 3.7 years.

Rolf Ent: Hall C

G0 completed their first commissioning run in G0 with 100 hours of `dry run' asymmetry data-taking. The results are encouraging, although there are concerns about the size of the inelastic background underneath the elastic proton peak, and differential non-linearity in the electronics and a large deadtime caused by the CFDs, both in the North American octants. The collaboration is concerned about the difficulties with delivering the 40 muA beam without disrupting other halls, but is eager to continue with the engineering run this fall.

From January to the end of March the Hall was switched over to the standard HMS/SOS configuration. The measurement of the F2 structure function at low Q2 was then performed, followed by the N-> Delta transition at high-Q2 experiment which is about half completed so far.

Changes to the vacuum beam dump line to reduce backgrounds were a big success. There are still problems with HMS controls systems; there a plans developed for an upgrade to fix these.

The schedule has the quark-hadron duality experiment at high-Q2 for 3 days, followed by the high-Q2 extension to the pion form factor experiment, then the experiment on quark-hadron duality in meson production experiment. Following the downtime will be the G0 engineering run starting mid-October, then a pause while Hall A runs the hypernuclear experiment, then the G0 forward-angle production run starting mid. February.

The HKS readiness review went well. Calorimeter construction for the GEp experiment is underway, and is about 1/2 done. This calorimeter is designed to be movable and can be used in experiments in Hall A as well as Hall C.

The Qweak experiment had a successful design review in January and has a completed project management plan. Funding has been secured from Canada (NSERC) and the NSF (via a 4-university MRI grant); negotiations for the DOE/JLab funding are underway.

The Hall C operations manual has undergone a major revision, and now consists primarily of a series of easy-to-follow `recipes' for procedures.

The backlog for experiments in Hall C is about 5 years.

User presentation

A member of the users group presented to the users group a situation involving a faculty member who has held a JLab Bridge position and who has been denied academic tenure by the university under what appear to be questionable justifications.

The Users group Board offered their support to the faculty member in question, and JLab management representatives also expressed their concern and consternation, and will investigate the matter with legal advisors etc.

Subsequently, a letter from the UGBOD Chair, representing the UBOD, expressing our strongest concern for what seems to us an arbitrary and unjust action, was sent to the Chancellor of the university involved.

As this is a personnel matter, and may involve legal proceedings, details are not reported here.

Space Issues: Pete Markowitz

Pete has had a chance to go over the drawings of the new space proposed in the CEBAF Center upgrade with Facilities Manager Rusty Sprouse. User input is welcome.

Computing Issues: Curtis Meyer

Graeme Heyes will be the new head of the Computer Center. There was concern expressed that this not mean a reduction in support/manpower for the DAQ group.

Running Experiments: Doug Higginbotham

The dipole in the injector used for energy measurement was moved (temporarily, we hope) for the Energy Recovery experiment. It was noted that a 5x10^-5 measurement of beam energy was made via the spin-dance measurement in 2000; Measurements of about 2x10^-4 precision can be made in Hall A with the combination of the Arc and e-p methods.

A new 499 MHz Ti-Sapphire laser (from the same vendor as the successful G0 laser) has been purchased; this laser has an excellent extinction ratio, and thus may be substantially reduce the beam leakage between the halls.

Quality of Life: Mark Pitt

There may be SURA money available for an upgrade of the computers in the residence facility. Mark will distribute a survey to determine how Users use the computers in the residence facility (eg. laptops with or without wireless cards? Do they depend on computers (eg thin clients) being provided in the rooms?).

A draft survey was distributed.

Workshop Document: Paul Stoler

After some discussion, it was decided not to attempt to produce formal proceedings of this week's Users Group workshop at the Annual Meeting. However, effort will be made to ensure that the transparencies will be made available on the WWW.

The meeting was adjourned at 10:15 PM